# Open-world Machine Learning: A Review and New Outlooks

### v2

@@ 4.2 Generalized Category Discovery
Generalize category discovery (GCD) is a more pragmatic and challenging task, relaxing the strong assumption in NCD that unlabeled data all come from novel classes, i.e., $\mathcal{Y}_{u}=\mathcal{C}_{\text{old}}\bigcap\mathcal{C}_{\text{new}}$ and $\mathcal{Y}_{l}\subset\mathcal{Y}_{u}$. The challenge in GCD lies in that models are supposed to simultaneously recognize unlabeled samples from old classes and cluster samples from novel categories. Illustrations of two canonical types of methods are shown in Figure 4.
@@
Non-parametric Classification. Most methods in GCD predominantly rely on contrastive learning for feature representation and resort to K-means [101, 102] clustering as a non-parametric classifier for classification. Vaze et al. [103] firstly introduced the task of GCD, and proposed to employ supervised contrastive learning [104] on $\mathcal{D}_{l}$ and unsupervised contrastive learning [54] on $\mathcal{D}_{u}$, respectively. Then semi-supervised K-means [101] is performed on the learned representations for classification of old samples and clustering of novel samples. They argue that parametric classifiers are prone to overfitting to old classes. This work could be further augmented by performing contrastive learning on sub-datasets partitioned by clustering in advance [105]. However, purely unsupervised contrastive learning neglects substantial positive pairs and suffers from the class collision issue, later works exploit more potential positive pairs to better feature representation. For example, a dynamic conceptional contrastive learning [106] framework is introduced to exploit the underlying relationships between samples of similar concepts. In this framework, dynamic conception generation and conception-level contrastive learning are alternatively performed for better feature representations. Zhao et al. [107] integrated class number estimation and representation learning into a unified EM-like framework to learn semi-supervised Gaussian mixture models. To further achieve better performance, PromptCAL [108] employs visual prompt tuning [109] with stronger representation ability.
Parametric Classification. Recently, Wen et al. [110] rethought parametric classifiers for GCD and attributed the failure of parametric classifiers to two biases, i.e., the bias to predicting old classes more often and the bias to imbalanced pseudo-labels. They proposed to implement self-distillation [111] for unlabeled data and employ entropy regularization to mitigate imbalanced pseudo-labels. By avoiding severe biases, the proposed method achieves remarkable performance with the parametric classifier.
4.3 Estimating the Number of Novel Classes
@@
In the literature of NCD and GCD, most methods require the number of novel classes $|\mathcal{C}_{\text{new}}|$ is known a-prior, which is impractical in real-world applications. It is of vital importance to precisely estimate the number of classes. AutoNovel [13] devise a sophisticated data splitting and clustering scheme. In particular, they first split the probe (labeled) data $\mathcal{D}_{l}^{r}$ into $\mathcal{D}_{l}^{ra}$ and $\mathcal{D}_{l}^{rv}$, and run semi-supervised K-means on $\mathcal{D}_{l}^{r}\bigcup\mathcal{D}_{u}$ with ground-truth constraints on $\mathcal{D}_{l}^{ra}$, subsequently, they proposed to compute clustering accuracy on $\mathcal{D}_{l}^{rv}$ and cluster validity index (CVI) [112] on $\mathcal{D}_{u}$, respectively. Finally, they average the cluster numbers maximizing accuracy and CVI respectively as the estimation of class number. Vaze et al. [103] proposed to run the clustering algorithm on the learned representations with different numbers of classes and choose the one that maximizes the accuracy on the labeled data as an estimate of $|\mathcal{C}_{\text{new}}|$. Recent methods adopt a more elegant way, which seamlessly integrates representation learning and category number estimation into a unified learning framework. DCCL [106] utilizes infomap algorithm [113] for dynamic conception generation, where the number of conceptions is dynamically changed throughout training, as the algorithm converges, and the resulting number of clusters represents the estimation. Zhao et al. [107] formalized EM-like framework, wherein prototype and class number are estimated in the E-step with a variant of Gaussian mixture model, and they employ prototypical contrastive learning [114] in the M-step.
4.4 Analysis of Category Discovery
@@
Novel class discovery is implicitly guided by knowledge transferred from old classes to novel categories for clustering and category discovery. How much knowledge could be transferred essentially influences the performance of category discovery. Li et al. [88] questioned the viewpoint that supervised knowledge is always helpful at different levels of semantic relevance. They proposed a metric called transfer flow to quantitatively assess the semantic similarity between labeled and unlabeled categories and give the theoretical analysis of semantic similarity and conclude that supervised knowledge might hurt the performance of NCD in circumstances with low semantic similarity between old and new classes and further suggest utilizing transfer flow as a reference to indicate how to use data.
@@
It is intuitive that the difficulty of NCD is determined by the semantic similarities between old and novel categories, i.e., how much knowledge is helpful and could be transferred for NCD. Besides, the number of novel classes also influences the difficulty. The greater the semantic similarity and the fewer the novel categories, the simpler the NCD task becomes. Interestingly, the impact of semantic similarity on the difficulty of novel class discovery is opposite to the conclusion of the unknown rejection task, in essence, unknown rejection is a discriminative task to distinguish between ID and OOD samples, while NCD is a transfer learning task to transfer the knowledge from ID to guide the clustering of OOD samples.
5 Class-incremental Learning
After detecting unknown samples via OOD detection or open-set recognition techniques, those samples should be labeled by humans or novel class discovery strategies. Then the system must continually extend the multi-class classifier to learn those new classes, which is referred to as class-incremental learning (CIL), being the third step in the open-world recognition process. Typically, an incremental learner learns several tasks sequentially, and only the data of the current task can be accessed by the learner. Most early studies [115, 116, 117] focus on task-incremental learning (TIL) scenario, in which the task identity is provided at inference time and the model only performs classification in each single task. Differently, the more challenging CIL problem involves the sequential learning of tasks that consist of disjoint class sets, and the model has to learn a unified classifier that can classify all seen classes. Formally, at incremental step $t$, a dataset $\mathcal{D}_{\text{train}}^{t}=\{(\bm{x}_{i}^{t},y_{i}^{t})\}^{N_{t}}_{i=1}$ is given, where $\bm{x}$ is an instance in the input space $\mathcal{X}$ and $y\in\mathcal{C}^{t}$ is its corresponding label. $\mathcal{C}^{t}$ is the class set of task $t$ and the class sets of different task are disjoint, i.e., $\mathcal{C}^{i}\cap\mathcal{C}^{j}=\emptyset$ if $i\neq j$. The general objective at each incremental step $t$ is to learn new classes $\mathcal{C}^{t}$ without interfering and with possibly improving previously learned knowledge [118].

### v3

@@ IV-C Generalized Category Discovery
Generalize category discovery (GCD) is a more pragmatic and challenging task, relaxing the strong assumption in NCD that unlabeled data all come from novel classes, i.e., $\mathcal{Y}_{u}=\mathcal{C}_{\text{old}}\bigcap\mathcal{C}_{\text{new}}$ and $\mathcal{Y}_{l}\subset\mathcal{Y}_{u}$. The challenge in GCD lies in that models are supposed to simultaneously recognize unlabeled samples from old classes and cluster samples from novel categories. Illustrations of two canonical types of methods are shown in Fig. 7.
1) Non-parametric Classification. Prior methods [72, 73, 75] in GCD predominantly rely on contrastive learning for feature representation. They do not explicitly learn a parametric classification head. Instead, they resort to K-means [323, 324] clustering as a non-parametric classifier for classification. Vaze et al. [72] firstly introduced the task of GCD, and proposed to employ supervised contrastive learning [325] on labeled data $\mathcal{D}_{l}$ (Eq. (7)) and unsupervised contrastive learning [268] on all data $\mathcal{D}$ (Eq. (8)), respectively.
$\mathcal{L}_{\text{con}}^{l}=\frac{1}{|\mathcal{B}_{l}|}\sum_{i\in\mathcal{B}_{l}}\frac{1}{|\mathcal{N}(i)|}\sum_{q\in\mathcal{N}(i)}-\log\frac{\exp(\mathbf{h}_{i}^{\top}\mathbf{h}_{q}/\tau_{c})}{\sum_{j}\mathds{1}_{[j\neq i]}\exp(\mathbf{h}_{i}^{\top}\mathbf{h}_{j}/\tau_{c})},$ (7)
$\mathcal{L}_{\text{con}}^{u}=\frac{1}{|\mathcal{B}|}\sum_{i\in\mathcal{B}}-\log\frac{\exp(\mathbf{h}_{i}^{\top}\mathbf{h}_{i}^{\prime}/\tau_{c})}{\sum_{j}\mathds{1}_{[j\neq i]}\exp(\mathbf{h}_{i}^{\top}\mathbf{h}_{j}/\tau_{c})}.$ (8)
@@
Here, $\mathcal{B}_{l}$ and $\mathcal{B}$ denote labeled samples and all samples in a batch, respectively. $\mathbf{h}$ represents the feature in projection space for contrastive learning. Then, semi-supervised K-means [323] is performed on the learned representations for the classification of old samples and clustering of novel samples. They argue that parametric classifiers are prone to overfitting to old classes. This work could be further augmented by performing contrastive learning on sub-datasets partitioned by clustering in advance [73]. However, purely unsupervised contrastive learning neglects substantial positive pairs and suffers from the class collision issue. Later works exploit more potential positive pairs to better feature representation. For example, a dynamic conceptual contrastive learning [74] framework is introduced to exploit the underlying relationships between samples of similar concepts. In this framework, dynamic conception generation and conception-level contrastive learning are alternatively performed for better feature representations. Zhao et al. [75] integrated class number estimation and representation learning into a unified EM-like framework to learn semi-supervised Gaussian mixture models. To further achieve better performance, PromptCAL [76] employs visual prompt tuning [326] with stronger representation ability from a diagonal perspective. InfoSieve [77] conceptualizes a category through the view of optimization and reformulates the task of category discovery as the assignment of minimum-length category codes for each sample. Hao et al. [78] take the spirit of agglomerative clustering [327] and construct a clustering hierarchy from the graph of selective neighbors.
2) Parametric Classification. Contrary to [72], Wen et al. [82] rethink parametric classifiers for GCD and attribute the failure of parametric classifiers to two biases, i.e., the bias to predicting old classes more often and the bias to imbalanced pseudo-labels. They utilize self-distillation [328] between two augmentation views for self-training, as shown below:
$\mathcal{L}_{\text{distill}}=\frac{1}{|\mathcal{B}|}\sum_{i\in\mathcal{B}}\ell(\bm{q}_{i}^{\prime},\bm{p}_{i}).$ (9)
Here $\ell(\cdot)$ is the cross-entropy loss and $\bm{q}_{i}^{\prime}$ is the soft target from the other view with a lower temperature, i.e., more confident distributions. Furthermore, entropy regularization is employed to mitigate imbalanced pseudo-labels and biased predictions toward old classes:
$\mathcal{L}_{\text{entropy}}=-H(\overline{\bm{p}})=\sum_{k=1}^{K}\overline{\bm{p}}^{(k)}\log\overline{\bm{p}}^{(k)},$ (10)
where $K$ is the total number of classes, and $H(\bm{p})=-\sum_{k}\bm{p}^{(k)}\log\bm{p^{(k)}}$ denotes entropy, and $\overline{\bm{p}}=\frac{1}{|\mathcal{B}|}\sum_{i\in\mathcal{B}}\bm{p}_{i}$ denotes marginal probability. By avoiding severe biases, the proposed method outperforms non-parametric methods by a large margin. $\mu$GCD [83] incorporates more advanced training schemes in FixMatch [329], e.g., exponential moving average and misaligned data augmentations to obtain more remarkable performance. SPTNet [84] proposes spatial prompt learning shared across images, empowering the model to focus on object parts. ProtoGCD [86] proposes a hybrid modeling with both discriminative and generative learning. Specifically, ProtoGCD models class-wise probability distribution with learnable prototypes. To balance the noise and self-learning efficiency of pseudo-labeling, dual-level adaptive pseudo-labels are introduced to consider the adaptivity across different training samples and phases. ProtoGCD also explicitly regularizes the class-wise prototypes to be more separable. Consequently, ProtoGCD achieves remarkable representations with large inter-class separation and intra-class compactness. The authors [86] also provide theoretical support for the design of pseudo-labels. Besides, PIM [81] provides a theoretical analysis for parametric methods from the perspective of mutual information. Generally, parametric classification methods are more efficient with end-to-end training. For inference, models could directly produce predictions with parametric classification heads without further clustering.
IV-D Estimating the Number of Novel Classes
@@
In the literature of NCD and GCD, most methods require the number of novel classes $|\mathcal{C}_{\text{new}}|$ to be known a priori, which is impractical in real-world applications. It is of vital importance to estimate the number of classes precisely. AutoNovel [58] devises a sophisticated data splitting and clustering scheme. In particular, they first split the probe (labeled) data $\mathcal{D}_{l}^{r}$ into $\mathcal{D}_{l}^{ra}$ and $\mathcal{D}_{l}^{rv}$, and run semi-supervised K-means on $\mathcal{D}_{l}^{r}\bigcup\mathcal{D}_{u}$ with ground-truth constraints on $\mathcal{D}_{l}^{ra}$, subsequently, they proposed to compute clustering accuracy on $\mathcal{D}_{l}^{rv}$ and cluster validity index (CVI) [330] on $\mathcal{D}_{u}$, respectively. Finally, they average the cluster number corresponding to the maximum accuracy and CVI value, as the estimation of the class number. Vaze et al. [72] proposed to run the clustering algorithm on the learned representations with different numbers of classes and choose the one that maximizes the accuracy on the labeled data as an estimate of $|\mathcal{C}_{\text{new}}|$. Recent methods adopt a more elegant way, which seamlessly integrates representation learning and category number estimation into a unified learning framework. DCCL [74] utilizes the infomap algorithm [331] for dynamic conception generation, where the number of conceptions is dynamically changed throughout training. As the algorithm converges, the resulting number of clusters represents the estimated results. Zhao et al. [75] formalized an EM-like framework, wherein prototype and class number are estimated in the E-step with a variant of the Gaussian mixture model, and they employ prototypical contrastive learning [332] in the M-step. A more recent work, ProtoGCD [86], devises prototype score, which simultaneously considers the accuracy of labeled data and the characteristics of feature space and achieves more precise estimation results.
IV-E Extended Settings
Recent efforts have extended GCD to other machine learning paradigms, like active learning [93] and federated learning [94]. Ma et al. [93] uncover the inherent imbalance problem between old and new classes in GCD. To effectively overcome this issue, they introduce the task of active generalized category discovery. By actively selecting valuable unlabeled samples, potentially from new classes and querying their labels from the oracle, the limited accuracy of new classes could be remarkably enhanced. To meet the privacy and decentralization requirements in practical scenarios, Pu et al. [94] formalize federated generalized category discovery. Furthermore, by considering the realistic streaming data, GCD could also be extended to the continual learning setting, i.e., continual category discovery [69, 68, 89, 88, 92], which could also be referred to as unsupervised continual learning. We will cover this content in Section V-F. Beyond the single visual modality, recent works [108, 109, 110] leverage orthogonal information from the text modality to further upgrade the performance of category discovery.
IV-F Theoretical Analysis of Novel Class Discovery
@@
Novelty class discovery is implicitly guided by knowledge transferred from old classes to novel categories for clustering and category discovery. How much knowledge could be transferred essentially influences the performance of category discovery. Li et al. [317] questioned the viewpoint that supervised knowledge is always helpful at different levels of semantic relevance. They proposed a metric called transfer flow to quantitatively assess the semantic similarity between labeled and unlabeled categories and give a theoretical analysis of semantic similarity and conclude that supervised knowledge might hurt the performance of NCD in circumstances with low semantic similarity between old and new classes. They further suggest utilizing transfer flow as a reference to indicate how to use the training data. Generally, the higher the semantic similarity between the old and new classes, the more substantial the contribution of the supervised training data to NCD.
As for the clustering mechanism of GCD, Chiaroni et al. [81] introduce constrained mutual information:
$\max_{\bm{W}}H(Y)-H(Y|Z),\quad\text{s.t.}\ \bm{y}_{i}=\bm{p}_{i},\quad\forall z_{i}\in Z_{l}.$ (11)
The overall objective is to maximize the marginal entropy $H(Y)$ while minimizing the conditional entropy of labels $Y$ conditioned on the underlying features $Z$. The first term corresponds to the equipartition assumption of class distributions. As for the second term, we need to assign the target $\bm{h}_{i}$ for each sample. In the context of GCD, there are both labeled data from old classes, as well as unlabeled samples from all classes. For the labeled samples, the annotations $\bm{y}_{i}$ are directly employed. For the unlabeled ones, they propose to use the predictive probabilities $\bm{p}_{i}$, which is equal to minimizing the predictive entropy of each sample, i.e., encouraging more confident predictions. In this way, the model performs semi-supervised learning and self-training.
@@
Intuitively, the difficulty of NCD is determined by the semantic similarities between old and novel categories, i.e., how much knowledge is helpful and could be transferred for NCD. Besides, the number of novel classes also influences the difficulty. The greater the semantic similarity and the fewer the novel categories, the simpler the NCD task becomes. Interestingly, the impact of semantic similarity on the difficulty of novel class discovery is opposite to the conclusion of the unknown rejection task, in essence, unknown rejection is a discriminative task to distinguish between ID and OOD samples, while NCD is a transfer learning task to transfer the knowledge from ID to guide the clustering of OOD samples.
TABLE IV: Comparative results of several NCD methods using task-aware evaluation protocol.
Method Venue CIFAR10 CIFAR100-20 CIFAR100-50 ImageNet
k-means [324] baseline 72.5 56.3 28.3 71.9
KCL [53] ICLR 2018 72.3 42.1 - 73.8
MCL [54] ICLR2018 70.9 21.5 - 74.4
DTC [55] ICCV 2019 88.7 67.3 35.9 78.3
RS [57] ICLR 2020 90.4 73.2 39.2 82.5
RS+ [58] TPAMI 2021 91.7 75.2 44.1 82.5
OpenMix [60] CVPR 2021 95.3 - - 85.7
[…]
DualRank [62] NeurIPS 2021 91.6 75.3 - 88.9
[…]
IIC [64] CVPR 2023 99.1 92.4 65.8 91.9
TABLE V: Comparative results of several NCD methods using task-agnostic evaluation protocol.
Method Venue CIFAR10 CIFAR100-20 CIFAR100-50
Labeled Unlabeled All Labeled Unlabeled All Labeled Unlabeled All
KCL [53] ICLR 2018 79.4 60.4 69.8 23.4 29.4 24.6 - - -
MCL [54] ICLR 2018 81.4 64.8 73.1 18.2 18.0 18.2 - - -
DTC [55] ICCV 2019 58.7 78.6 68.7 47.6 49.1 47.9 30.2 34.7 32.5
RS+ [58] TPAMI 2021 90.6 88.8 89.7 71.2 56.8 68.3 69.7 40.9 55.3
[…]
TABLE VI: Comparative results of several GCD methods on generic datasets.
Type Methods Venue CIFAR10 CIFAR100 ImageNet-100
All Old New All Old New All Old New
Non-parametric Classification K-means [324] Baseline 83.6 85.7 82.5 52.0 52.2 50.8 72.7 75.5 71.3
GCD [72] CVPR 2022 91.5 97.9 88.2 73.0 76.2 66.5 74.1 89.8 66.3
XCon [73] BMVC 2022 96.0 97.3 95.4 74.2 81.2 60.3 77.6 93.5 69.7
DCCL [74] CVPR 2023 96.3 96.5 96.9 75.3 76.8 70.2 80.5 90.5 76.2
[…]
InfoSieve [77] NeurIPS 2023 94.8 97.7 93.4 78.3 82.2 70.5 80.5 93.8 73.8
[…]
Parametric Classification RankStats+ [58] TPAMI 2021 46.8 19.2 60.5 58.2 77.6 19.3 37.1 61.6 24.8
UNO+ [61] ICCV 2021 68.6 98.3 53.8 69.5 80.6 47.2 70.3 95.0 57.9
ORCA [80] ICLR 2022 81.8 86.2 79.6 69.0 77.4 52.0 73.5 92.6 63.9
[…]
SimGCD [82] ICCV 2023 97.1 95.1 98.1 80.1 81.2 77.8 83.0 93.1 77.9
SPTNet [84] ICLR 2024 97.3 95.0 98.6 81.3 84.3 75.6 85.4 93.2 81.4
[…]
TABLE VII: Comparative results of several GCD methods on fine-grained datasets.
Type Methods Venue CUB Stanford Cars FGVC-Aircraft Herbarium19
All Old New All Old New All Old New All Old New
Non-parametric Classification K-means [324] Baseline 34.3 38.9 32.1 12.8 10.6 13.8 16.0 14.4 16.8 13.0 12.2 13.4
GCD [72] CVPR 2022 51.3 56.6 48.7 39.0 57.6 29.9 45.0 41.1 46.9 35.4 51.0 27.0
XCon [73] BMVC 2022 52.1 54.3 51.0 10.5 58.8 31.7 47.7 44.4 49.4 - - -
DCCL [74] CVPR 2023 63.5 60.8 64.9 43.1 55.7 36.2 - - - - - -
GPC [75] ICCV 2023 55.4 58.2 53.1 42.8 59.2 32.8 46.3 42.5 47.9 36.5 51.7 27.9
PromptCAL [76] CVPR 2023 62.9 64.4 62.1 50.2 70.1 40.6 52.2 52.2 52.3 37.0 52.0 28.9
InfoSieve [77] NeurIPS 2023 69.4 77.9 65.2 55.7 74.8 46.4 56.3 63.7 62.5 41.0 55.4 33.2
CiPR [78] TMLR 2024 57.1 58.7 55.6 47.0 61.5 40.1 - - - 36.8 45.4 32.6
[…]
Parametric Classification RankStats+ [58] TPAMI 2021 33.3 51.6 24.2 28.3 61.8 12.1 26.9 36.4 22.2 27.4 55.8 12.8
UNO+ [61] ICCV 2021 35.1 49.0 28.1 35.5 70.5 18.6 40.3 56.4 32.2 28.3 53.7 14.7
ORCA [80] ICLR 2022 35.3 45.6 30.2 23.5 50.1 10.7 22.0 31.8 17.1 20.9 30.9 15.5
PIM [81] ICCV 2023 62.7 75.7 56.2 43.1 66.9 31.6 - - - 42.3 56.1 34.8
[…]
$\mu$GCD [83] NeurIPS 2023 65.7 68.0 64.6 56.5 68.1 50.9 53.8 55.4 53.0 - - -
[…]
IV-G Evaluation of Novel Class Discovery
1) Evaluation of NCD. Canonically, NCD is evaluated using two protocols: (1) task-aware, where each test sample is known to be from old or novel classes in advance, and (2) task-agnostic, where the prior information is not given and the model needs to classify the sample from all classes. The classification accuracy and cluster accuracy (see Eq. (12) for more details) are reported for old and new classes, respectively. The results of task-aware and task-agnostic are elaborated in Table IV and Table V, respectively. By default, all methods use ResNet-18 [271] as the backbone. All methods are evaluated on CIFAR10 [311], CIFAR100 [311] and ImageNet [333] datasets. Generally, more recent methods ComEx [63] and IIC [64] achieve remarkable performance.
2) Evaluation of GCD. In essence, GCD is a clustering task for both old and new categories. One could not directly compare model predictions and ground truth labels like in conventional supervised learning. In GCD, the clustering accuracy (ACC) of the model’s predictions $\tilde{y}_{i}$ given the ground-truth labels $y_{i}$ is measured as:
$ACC=\max_{\omega\in\Omega(\mathcal{Y}_{u})}\frac{1}{M}\sum_{i=1}^{M}\mathds{1}\big{\{}y_{i}=\omega(\tilde{y}_{i})\big{\}}.$ (12)
Here, $\mathds{1}(\cdot)$ is the indicator function, and $M$ denotes the number of test samples. $\Omega(\mathcal{Y}_{u})$ represents the set of all bijective permutations that match the model predictions to the ground-truth labels. The optimal permutation in Eq. (12) is obtained through the Hungarian algorithm [334], which is only implemented once across all classes. Here, we present a comprehensive comparative evaluation of several GCD methods, including both non-parametric and parametric types. The results on generic and fine-grained datasets are elaborated in Table VI and Table VII, respectively. By default, the backbone selects DINO [328] pre-trained ViT-B/16 [272], and only the last transformer block is tuned. The [CLS] token of the last layer is selected as the feature representation. The evaluation datasets include generic image classification datasets in Table VI, e.g., CIFAR10 [311], CIFAR100 [311] and ImageNet [333] subsets, as well as fine-grained datasets in Table VII, like CUB [335], Stanford Cars [336], FGVC-Aircraft [337] and Herbarium 19 [338]. By default, 50% of the classes serve as the old classes, while others serve as novel classes. Overall, the recent parametric classification methods [83, 84, 86] achieve better results.
Figure 8: (a) Updating the model continuously on new classes without accessing old data leads to three phenomena: (b) (old-old) representation and classifier mismatch, (c) (old-new) representation confusion, and (d) (old-new) classifier imbalance. Those three issues result in the catastrophic forgetting problem.
V Continual Learning
After detecting unknown samples via OOD detection or open-set recognition techniques (Sec. III), those samples should be labeled by humans or novel class discovery strategies (Sec. IV). Then, the system must continually extend the multi-class classifier to learn those new classes, which is referred to as continual learning (CL), being the third step in the open-world recognition process (Fig. 2). In this section, we provide a comprehensive survey on recent advances in CL. To be specific, we first present the problem definition and key challenges in Sec. V-A. Then, we group existing CL studies into several main categories: regularization methods (Sec. V-B), data replay methods (Sec. V-C), feature replay methods (Sec. V-D) and model expansion methods (Sec. V-E). Then, we demonstrate some recent extended practical settings of CL (Sec. V-F) and theoretical advance of continual learning (Sec. V-G). Finally, some evaluation results are presented (Sec. V-H).
V-A Problem Formulation and Key Challenges
1) Problem Formulation. Typically, a CL problem involves the sequential learning of tasks that consist of disjoint class sets, and the model has to learn a unified classifier that can classify all seen classes. At incremental step $t$, a dataset $\mathcal{D}_{t}=\{\bm{x}_{i}^{t},y_{i}^{t}\}^{n_{t}}_{i=1}$ is given, where $\bm{x}$ is an image in the input space $\mathcal{X}$ and $y\in\mathcal{C}_{t}$ is its corresponding label. $\mathcal{C}_{t}$ is the class set of task $t$ and the class sets of different task are disjoint, i.e., $\mathcal{C}_{i}\cap\mathcal{C}_{j}=\emptyset$ if $i\neq j$. To facilitate analysis, we represent the DNN-based model with two components: a feature extractor and a unified classifier. Specifically, the feature extractor $f_{\bm{\theta}}:\mathcal{X}\rightarrow\mathcal{Z}$, parameterized by $\bm{\theta}$, maps the input $\bm{x}$ into a feature vector $\bm{z}=f_{\bm{\theta}}(\bm{x})\in\mathbb{R}^{d}$ in the feature space $\mathcal{Z}$; the classifier $g_{\bm{\varphi}}:\mathcal{Z}\rightarrow\mathbb{R}^{|\mathcal{C}_{1:t}|}$, parameterized by $\bm{\varphi}$, produces a probability distribution $g_{\bm{\varphi}}(\bm{z})$ as the prediction for $\bm{x}$. Denote the overall parameters by $\bm{\Theta}=(\bm{\theta},\bm{\varphi})$. At stage $t$, the general objective is to minimize a predefined loss function $\ell$ (i.e., cross-entropy loss) on the new dataset $\mathcal{D}_{t}$ without interfering and with possibly improving on those that were learned previously [339]:
$\operatorname*{argmin}\limits_{\bm{\theta},\bm{\varphi},\epsilon}\mathbb{E}_{(\bm{x},y)\backsim\mathcal{D}_{t}}[\ell($ $g_{\bm{\varphi}}(f_{\bm{\theta}}(\bm{x})),y)]+\sum\nolimits\epsilon_{i}$ (13)
[…]
$\epsilon_{i}\geqslant 0;~{}~{}\forall$ $i\in[1,t-1].$
The last term $\epsilon=\{\epsilon_{i}\}$ is a slack variable that tolerates a small increase in the loss on datasets of old tasks.
2) Key Challenges. The central challenge of CL is that old data are unavailable when learning new classes and the objective defined in Eq. 13 can not be optimized directly. Therefore, the best configuration of the model for all seen classes must be sought by minimizing $\mathcal{L}_{t}$ on $\mathcal{D}_{t}$:
$\mathcal{L}_{t}=\mathbb{E}_{(\bm{x},y)\backsim\mathcal{D}_{t}}[\ell(g_{\bm{\varphi}}(f_{\bm{\theta}}(\bm{x})),y)].$ (14)